Subject: belated) APO 3.5m users committee minutes, June 2011
From: Suzanne Hawley
Submitted: Wed, 13 Jul 2011 15:11:22 -0700 (PDT)
Message number: 1233
(previous: 1232,
next: 1234
up: Index)
NOTE: Information item on echelle received after the June meeting
The ThAr lamp was replaced on June 6, and the replacement lamp has a
quartz window which makes many of the lines, especially in the blue, very
faint and unusable for wavelength calibration. Echelle users who have data
from before June 6 are warned that they will likely not be able to use
their existing wavecal template. Instead, many of the faint blue lines
must be manually removed from the wavelength solution. We will discuss
further at the next users committee meeting on August 8.
***********************************
APO 3.5-m Users Committee Phonecon, 6/06/11
Attending: Suzanne Hawley, Al Harper, Bill Ketzeback, Bruce Gillespie,
John Bally, Jon Holtzman, Michael Strauss, Russet McMillan, Scott Anderson
Absent: Mark Klaene, Remy Indebetouw, Sean Moran
**********************************
User feedback and comments from institutional representatives:
o Washington (Scott Anderson) - Scott had nothing new to report.
o Colorado (John Bally) - John had nothing new to report.
o Johns Hopkins (Sean Moran) - No report.
o New Mexico State (Jon Holtzman) - Jon had nothing new to report.
o Princeton (Michael Strauss) - Michael had nothing new to report.
o Virginia (Remy Indebetouw) - No report.
o Chicago (Al Harper) - Al had nothing new to report.
**********************************
Discussion of telescope/instruments report:
Bill Ketzeback gave an overview of the past month's activities. Although
forest fires have been and are burning in the local area, there has been
no significant impact to site operations. We have had a small amount of
rain in the past week, the first since January, and although the rain is
appreciated we worry about so-called dry lightning. Observing conditions
have been generally good. The telescope has behaved well; a small change
in implementing pointing model corrections was effected that should
improve telescope pointing and tracking slightly. The replacement camera
for the Shack-Hartmann wavefront sensor has just arrived, and will be
installed in the coming weeks. The driver card for the DIS grating turret
motor failed, and has been fixed in spite of trouble using the
documentation for it. SPIcam is OK, although it is running a little
warmer than we would like. The problem with the Agile instrument rotator
was fixed. Some shims and adjustment to the NA2 clamps has fixed the
TripleSpec mounting problem, but a generic redesign of the clamps in being
considered. The echelle shutter failed this past weekend [ed: since fixed
by wiggling a cable, which will be replaced]. The NIC-FPS reset failures
when upside down persists -- the operational workaround is being used, and
the instrument will go back to CU to have its internal cables replaced
next month. The Goddard team is back at the site to continue the
integral-field and detector upgrades to the GFP.
Mark Klaene's detailed report for the reporting period follows:
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights
5/6/11 through 6/1/11
0) Overview
The forest closed on May 12 due to critical fire danger. Immediately
following, a large fire, caused by a downed power line, started 20 miles
northeast of us, ultimately burning 32,000 acres. APO operations have not
been affected by the closure. Despite being through May, the winds have
continued near the levels of April. Short telescope closures due to dust
and smoke have occurred a few times during this period.
1) Telescope
The telescope continues to operate routinely with minimal problems during
this period. Some minor changes to pointing models have been
incorporated.
The Shack-Hartmann camera controller failed and a new camera and
controller have been purchased. We will be replacing the Apogee camera
with a new SBIG camera. Camera delivery is expected this month.
2) Instruments
DIS performed well this period until a failure of the turret motion. The
instrument was configured that the turret could be moved by hand which
causes a loss of about 15 minutes if grating sets need to be changed
during the night. The cryotiger was adjusted but it seems to have had
little effect. The camera is still running 10-12 deg C warmer than
previous when ambient temperatures are warmer. We hope to wait until
summer shutdown to investigate further.
SPICam operated with no new issues.
Agile rotator fault was corrected; a problem induced by the new locking
pin was found to be the cause.
A TripleSpec problem has been diagnosed where the instrument comes off the
kinematic mounts slightly when rotated upside down. No effect on science
has been reported and the instrument safety is not at risk. A temporary
solution to adjust the clamps and shim the pawls on the instrument was
done and appears to have been successful. New clamping/mounting systems
are being investigated.
The Echelle worked with no problems this period.
The NIC-FPS problem of spontaneous reset of the Leach controller when
rotated upside down still exists. Users are advised to start their
observations at specific recommended rotation angles, depending on sky
position, to avoid tracking upside down. The instrument will be taken to
CU on June 20 for a period of approximately 6 weeks.
GFP upgrade work has resumed on the mountain with an engineering run
coming up next week.
**********************************
Reporting JD in FITS headers:
Suzanne said that there is a request to report Julian date in the FITS
image headers. The issue is which time system to use as the standard.
SDSS reports MJD in TA1, which is what we are suggesting for the 3.5-m
instruments. Michael felt that this was fine, if things are
well-labelled. Suzanne suggested that we can describe this on the web in
the user documentation. Bill said that the instrument computers can and
do report in multiple time systems, especially the older computers. So
setting a MJD-in-TA1 standard would require changing some software for the
older instruments. It was generally felt that having a standard time
system was probably a good idea, but that we should see if the users
really want it to justify the extra work involved.
===> ACTION: Users Committee members to ask their users if MJD in TA1 would
be useful information to have written in the FITS headers.
**********************************
Possible new instrument(s):
Suzanne mentioned a couple of items. First, we discussed several months
ago the possibility of getting a multi-band griz imager that Chris Stubbs
is involved with at CfA, called PISCO. Chris in interested in possibly
deploying the camera at APO and will send us a proposal. The terms of its
use at the 3.5-m telescope would need to be negotiated. The camera has a
FOV of approx. 10 arcminutes. There are questions about telescope
baffling changes, and whether the filters are changeable. We need more
feedback on the desirability of PISCO from the user committee and general
users. Michael mentioned the Princeton would be very interested in the
instrument. Jon Holtzman was concerned that PISCO doesn't have all the
capabilities that SPIcam has (i.e., narrow-band filters), so if PISCO
replaced SPIcam, it is not clear that this is win-win. There is concern
that SPIcam will fail in the near future and not be easily repaired.
There is also a floor-space issue on the NA2 side of the telescope -- we
may need to retire an instrument in order to find room for a new one. In
any case, we need to hear back from the user community about the pros and
cons of this possible new instrument.
Second, Jon Holtzman has been talking with Matthew Shetrone at UT about
IFUs. The HET is having several built, and it may be possible to get one
additional copy made for use at APO. It's not clear if this IFU would be
better than the IFU upgrade underway for the GFP. Michael said that there
may be people at Princeton interested in an IFU on the 3.5-m.
**********************************
Misc. topic - ReSTAR:
Jon mentioned that he got a note from NOAO that they are starting to write
the joint ReSTAR proposal for NSF funding, and we need to decide on the
terms of our participation. Discussions are ongoing.
**********************************
Summer telecon schedule:
Suzanne said that there will be no Users Committee telecon in July during
the summer shutdown. The next telecon will be 8 August. The following
telecon will be in late September or early October.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: User committee members to try to find users who would be
interested in helping develop exposure time calculators.
STATUS: Scott said that Anjum Mukadam is willing to develop exposure
calculators for Agile, and likewise Sarah Schmidt for TripleSpec. Suzanne
said that George Wallerstein is willing to help with the echelle. Jon noted
that this really needs to be done in a single framework, and Bruce suggested
that we import an existing exposure calculator architecture from NOAO or
STScI, for example. Suzanne decided to leave the action item open, giving
the Users Committee members more time to ask for help and to look into the
possibility of importing a framework from elsewhere.
[new actions from this meeting]:
===> ACTION: User Committee members to ask their users if MJD in TA1 would be
useful information to have written in the FITS headers.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 8 August
at 8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1233 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO